OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (alexandervonhumboldt2):

Focus: (2, 6) Any point, (x0 , y0) on the parabola satisfies the definition of parabola, so there are two distances to calculate: Distance between the point on the parabola to the focus Distance between the point on the parabola to the directrix \[\sqrt{(x_0-a)^2+(y_0-b)^2} \] \[\sqrt{(x_0-2)^2+(y_0-0)^2}\] y=-2 dirextix Distance between point ( x0 , y0) and the line y=-2 |y_0+2| Equate the two expressions. \[\sqrt{(x_0-2)^2+(y_0-0)^2}=\left| y_0+2 \right|\] now solve this:

OpenStudy (alexandervonhumboldt2):

square both sides: \[(x_0-2)^2+y_0^2=(y_0+2)^2\]

OpenStudy (alexandervonhumboldt2):

simplify: \[x^2-2x+4+y^2=y^2+2y+4\]

OpenStudy (alexandervonhumboldt2):

simplify: \[x^2-2x+4-4+y^2-y^2=2y\]

OpenStudy (alexandervonhumboldt2):

\[x^2-2x=2y\]

OpenStudy (alexandervonhumboldt2):

\[y=\frac{ x^2-2x }{ 2 }\]

OpenStudy (alexandervonhumboldt2):

i might have done a mistake though haven't made such question for long time

OpenStudy (anonymous):

thats not any of my choices :/

OpenStudy (alexandervonhumboldt2):

what are your choises?

OpenStudy (anonymous):

\[f(x)= -\frac{ 1 }{ 8 } (x-2)^{2}-2\]

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 16 } (x-2)^{2}-2\]

OpenStudy (anonymous):

and then the same as the first except positive 1/8 and same as the second but negative 1/16

OpenStudy (alexandervonhumboldt2):

OHHHHHHHHHHHHH i counted with focus (2, 0) not (2, 6) wait lemme redo it

OpenStudy (anonymous):

okey

OpenStudy (alexandervonhumboldt2):

\[\sqrt{(x_0-2)^2+(y_0-6)^2}=\left| y_0+2 \right|\]

OpenStudy (alexandervonhumboldt2):

\[(x_0-2)^2+(y_0-6)^2=(y_0+2)^2\]

OpenStudy (alexandervonhumboldt2):

\[(x-2)^2+y^2-12y+36=y^2+2y+4\]

OpenStudy (alexandervonhumboldt2):

\[(x-2)^2+32=14y\]

OpenStudy (alexandervonhumboldt2):

woops mistake again

OpenStudy (alexandervonhumboldt2):

(x-2)^2+y^2-12y+36=y^2+4y+4

OpenStudy (alexandervonhumboldt2):

(x-2)^2+32=16y y=1/16*(x-2)^2+2 i the answer if i again didn't made a mistake somewhere

OpenStudy (anonymous):

thanks! ill let you know in 1 sec

OpenStudy (anonymous):

yep! it was right!! i got 100/100 on my quiz thanks!

OpenStudy (alexandervonhumboldt2):

np sorry for taking so much time

OpenStudy (alexandervonhumboldt2):

np sorry for taking so much time